OpenStudy (anonymous):

Tammy is at the dentist's office waiting on her appointment. She notices that the 6-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 1:20 to 1:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? You must show all of your work.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Were you able to solve the hint? Degrees per minute = ?

OpenStudy (anonymous):

No. I wasn't :(

OpenStudy (anonymous):

was it 210?

zepdrix (zepdrix):

A full rotation is 360 degrees. A full rotation is also 60 minutes, yes? So what is the ratio of degrees to minutes?\[\large\rm \frac{360}{60}=?\]

OpenStudy (anonymous):

6

zepdrix (zepdrix):

So there are 6 degrees in each minute. Ok great.

zepdrix (zepdrix):

So as follow up questions I'll ask: `How many minutes are there between 1:20 and 1:55?` and then using the Hint, `How many degrees are there between 1:20 and 1:55?`

OpenStudy (anonymous):

hmm so i am assuming 210 wouldnt be the answer right?

OpenStudy (anonymous):

oh okay sorrry! 35 mintues

zepdrix (zepdrix):

Ok great so you've answered my second question. 210 degrees. Remember how to convert to radians? :)

OpenStudy (anonymous):

um is it 180 times x/pi?

OpenStudy (anonymous):

is it 35 min. from 1:20-1:55

zepdrix (zepdrix):

35 minutes from 1:20 to 1:55 6 degrees per minute, 35*6 degrees from 1:20 to 1:55. Yes, good. Hopefully that part makes sense.

zepdrix (zepdrix):

180 degrees and pi are equivalent on our unit circle.\[\large\rm 180^o=\pi\]We can either divide by pi to get,\[\large\rm \color{royalblue}{\frac{180^o}{\pi}=1}\]Or we can instead divide by 180 degrees to get,\[\large\rm \color{royalblue}{\frac{\pi}{180^o}=1}\]We want to use one of these blue relationships to deal with our 210 degrees.

OpenStudy (anonymous):

Yay! Ok I see

zepdrix (zepdrix):

You have 210 degrees. Think of that value like this: \(\large\rm 210^o=\dfrac{210^o}{1}\) So if our degrees are in the numerator, which fraction should we use to cancel out the degrees and end up with radians?

zepdrix (zepdrix):

Which blue one?

OpenStudy (anonymous):

2nd one?

zepdrix (zepdrix):

Ahh, well put detective! \[\large\rm \frac{210^o}{1}\cdot\color{royalblue}{1}\quad=\frac{210^{o}}{1}\cdot\color{royalblue}{\frac{\pi}{180^{o}}}\]Degrees in the bottom will cancel with our degrees in the top,\[\large\rm \frac{210^{\cancel o}}{1}\cdot\frac{\pi}{180^{\cancel o}}\quad=\frac{210\pi}{180}\]and simplify.

OpenStudy (anonymous):

hmm ok I'm bad at simplifying fractions, so I'll go with 7pi/12?

zepdrix (zepdrix):

So taking the zero out is a nice first step. 21pi/18 Then I think we can take out.... a 3 from that point, ya?

OpenStudy (anonymous):

yes:)

zepdrix (zepdrix):

Hmm 18 divided by 3 is not 12. :d

OpenStudy (anonymous):

it is 6

zepdrix (zepdrix):

Oh I guess I made a little boo boo. Yes, we have a radian measure of 7pi/6, good. But that's for a circle of radius 1. Notice that our minute hand is 6 inches, so our radius is 6.

OpenStudy (anonymous):

ok so the answer to part 1 would be 7pi/6?

zepdrix (zepdrix):

No, we'll need to multiply that value by 6 to correspond to a circle of radius 6 :)

OpenStudy (anonymous):

Oh yuck! Ok let's do it:)

zepdrix (zepdrix):

It's not too bad. You have a fraction being divided by 6, and you're multiplying that by 6. So we just lose our denominator, ya?

OpenStudy (anonymous):

yup!

zepdrix (zepdrix):

So for part 1, I suppose we end up with 7pi. Ok not too bad :o

zepdrix (zepdrix):

Oh maybe I confused parts 1 and 2.. hmm thinking... because now they're asking about the `tip` of the minute hand... mmm

zepdrix (zepdrix):

Ooo boy, this is a poorly worded question :( But `I think` what they want is 7pi/6 for Part 1, and 7pi for part 2, we multiply the radian measure by the radius to get the arc length.

zepdrix (zepdrix):

Mmm what do you think? You as confused as I am? 0_o

OpenStudy (anonymous):

Answer is 10pi/3:)